OpenStudy (anonymous):

A 450g cube of gold at 225 degrees Celsius is cooled by dunking it in 850g of water at 20 degrees C. what is the final temperature of the mixture? specific heat of gold = 0.031 cal/gC specific heat of water = 1

OpenStudy (anonymous):

q lost = q gain So, by substitution, we then have: (450) (225 - x)(0.031) = (850) (x - 20) (1.0) Solve for x 13.95 (225 - x) = 850 (x - 20) 3138.75 - 13.95x = 850x - 17000 then 863.95x = 20138.75 The answer is 23.31 °C

OpenStudy (shamim):

U know heat loss by gold is just equal to the heat gained by water

OpenStudy (shamim):

We assume this water and gold r isolated frm environment

OpenStudy (shamim):

So no heat transfer will occur between the system ( water n gold ) and environment

OpenStudy (shamim):

Here used Q=ms*del theta

OpenStudy (shamim):

Q= heat loss or heat gain

OpenStudy (shamim):

m=mass of object

OpenStudy (shamim):

S=specific heat of the object

OpenStudy (shamim):

Del theta= temperature difference